# Bertrand’s and Rodriguez Villegas’ Conjecture for real quartic Galois extensions of the rationals

### v1

@@ 4. The Cyclic Group Case
In this section, we prove the main theorem for the case when  $\operatorname{Gal}(L/\mathbb{Q})$ is isomorphic to the cyclic group. Throughout this section, let $L$ be a real Galois extension of $\mathbb{Q}$ with $\operatorname{Gal}(L/\mathbb{Q})\cong\mathbb{Z}/4\mathbb{Z}$. Denote the Galois group as  $G$. Let $\operatorname{Gal}(L/\mathbb{Q})$ be generated by $\sigma$. Let $l$ be the unique quadratic subfield of $L$. See Figure 2. Let $u_{l}>1$ be the fundamental unit of $l$. Then ${\mathcal{O}_{l}^{*}}=\{\pm u_{l}^{m}\ |\ m\in\mathbb{Z}\}$. First we define relative units as in [5].
$\mathbb{Q}$$l$$L$$\{1,\sigma^{2}\}$$G=\{1,\sigma,\sigma^{2},\sigma^{3}\}$$\{1\}$
Figure 2. Subfields of cyclic extension $L$ and corresponding Galois groups
@@
Let $E_{L/l}$ be a group consisting of all relative units of $L$. As in [4], let $E^{L}$ be a sub $G$-module of ${\mathcal{O}_{L}^{*}}$ generated by elements of  ${\mathcal{O}_{l}^{*}}$ and $E_{L/l}$. Let $Q=[{\mathcal{O}_{L}^{*}}:E^{L}]$. The following lemmas are from Hasse [5].
@@
There exists an element $u_{0}$ of $E_{L/l}$ such that $-1,u_{0}$ and its conjugate $\sigma(u_{0})$ generates $E_{L/l}$.
Proof.
See Satz 22, p.40 of Hasse [5]. ∎
@@
where $W_{1}=\log(u_{l}),W_{2}=\log(\lvert u_{0}\rvert),W_{3}=\log(\lvert\sigma(u_{0})\rvert)$. Here, the order of the basis of $\mathbb{R}^{\mathcal{A}_{L}}$ is $\delta^{id},\delta^{\sigma},\delta^{\sigma^{2}},\delta^{\sigma^{3}}$. Note that $W_{1},W_{2},W_{3}$ are nonzero since $u_{l},u_{0},\sigma(u_{0})$ are not ±$\pm 1$. Then the wedge product of these terms are:
@@
(4.2) $f(n_{1},n_{2},n_{3})\geq 4\lvert n_{3}Y_{3}\rvert=4\lvert n_{3}\rvert(W_{2}^{2}+W_{3}^{2})$
@@
(4.3) $f(n_{1},n_{2},n_{3})\geq 2W_{1}(2\max\{\lvert W_{2}\rvert,\lvert W_{3}\rvert\}+\lvert W_{2}\rvert+\lvert W_{3}\rvert)$
Using these results, we prove the Theorem 1.2 for each case $Q=1$ and $Q=2$.
@@
Let $X,Y\in\mathbb{R}$. For any $(m,n)\in\mathbb{Z}^{2}\backslash\{(0,0)\}$, the following inequality
$\lvert mX+nY\rvert+\lvert nX-mY\rvert\geq\lvert X\rvert+\lvert Y\rvert$
holds.
@@
$\lvert mX+nY\rvert+\lvert nX-mY\rvert=\lvert X\rvert(\lvert m+nZ\rvert+\lvert n-mZ\rvert),$
@@
Without loss of generality, assume $m\geq 0$. If $m=0$$\lvert nZ\rvert+\lvert n\rvert\geq 1+Z$ for any nonzero integer $n$. If $m\geq 2$,
$\lvert m+nZ\rvert+\lvert n-mZ\rvert$ $\geq\lvert m+nZ\rvert+\lvert Z(n-mZ)\rvert$
$\geq\lvert(m+nZ)-Z(n-mZ)\rvert$
$=\lvert m(1+Z^{2})\rvert$
$\geq m$
$\geq 2$
$\geq 1+Z.$
@@
$\lvert 1+nZ\rvert+\lvert n-Z\rvert\geq{n-Z}\geq 1+Z$
@@ 4.1. The Case Q=1
$\lVert w\rVert_{1}\geq 2W_{1}(2\max\{\lvert W_{2}\rvert,\lvert W_{3}\rvert\}+\lvert W_{2}\rvert+\lvert W_{3}\rvert).$
@@
(4.4) $\lVert w\rVert_{1}\geq\min\{4(W_{2}^{2}+W_{3}^{2}),2W_{1}(2\max\{\lvert W_{2}\rvert,\lvert W_{3}\rvert\}+\lvert W_{2}\rvert+\lvert W_{3}\rvert)\}$
@@
Let $L$ be totally real number field and $u\in{\mathcal{O}_{L}^{*}}$ with $u\neq\pm 1$. Then
$\lVert\operatorname{LOG}(u)\rVert_{2}^{2}\geq[L:\mathbb{Q}]\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}$
holds.
@@
Bemerkung in p.102 of Pohst [7] states that the Lemma in p.98 applies for  $M$ of the eqaution (10) in p.97. Then the inequality (11) of the Lemma gives the required result. ∎
@@
(4.5) $W_{2}^{2}+W_{3}^{2}=\frac{1}{2}\lVert\operatorname{LOG}(u_{0})\rVert_{2}^{2}\geq 2\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}.$
Under this constraint and the fact that $W_{2},W_{3}$ are nonzero, an elementary consideration yields
$2\max\{\lvert W_{2}\rvert,\lvert W_{3}\rvert\}+\lvert W_{2}\rvert+\lvert W_{3}\rvert>3\sqrt{2}\log\left(\frac{1+\sqrt{5}}{2}\right).$
@@
(4.6) $W_{1}=\log(u_{l})\geq\log\left(\frac{1+\sqrt{5}}{2}\right)$
holds, and with previous inequality,
$2W_{1}(2\max\{\lvert W_{2}\rvert,\lvert W_{3}\rvert\}+\lvert W_{2}\rvert+\lvert W_{3}\rvert)>6\sqrt{2}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}.$
@@
(4.7) $\lVert w\rVert_{1}\geq\min\left\{8\log\left(\frac{1+\sqrt{5}}{2}\right)^{2},6\sqrt{2}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}\right\}=8\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}$
@@ 4.2. The Case Q=2
We use the following lemma from Hasse [5].
@@
If $Q=2$, there exists a unique unit $u_{*}>0$ of $L$ that satisfy
(4.8) $N_{L/l}(u_{*})=u_{*}\sigma^{2}(u_{*})=\pm u_{l},\quad u_{*}\sigma(u_{*})=\pm u_{0}.$
Proof.
See Satz 28, p.45 of Hasse [5]. ∎
@@
and therefore  $\sigma(u_{0})=\pm u_{l}u_{0}u_{*}^{-2}$. The proof of Lemma 4.2 shows that when $Q=2$, every unit $u\in{\mathcal{O}_{L}^{*}}$ can be written uniquely as the form ±$\pm u_{*}^{n_{0}}u_{l}^{n_{1}}u_{0}^{n_{2}}\sigma(u_{0})^{n_{3}}$ where $n_{0}$ is 00 or $1$ and $n_{1},n_{2},n_{3}$ are integers. Thus  ${\mathcal{O}_{L}^{*}}$ is generated by $-1,u_{l},u_{0},u_{*}$. Hence  $\bigwedge^{2}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$ is generated by $\operatorname{LOG}(u_{l})\wedge\operatorname{LOG}(u_{0}),\operatorname{LOG}(u_{l})\wedge\operatorname{LOG}(u_{*})$ and  $\operatorname{LOG}(u_{0})\wedge\operatorname{LOG}(u_{*})$. From (4.9), we also have
@@
We now proceed to give a lower bound for $f(n_{1},n_{2},n_{3})$ where $n_{1}+n_{2}+n_{3}$ is even and $(n_{1},n_{2},n_{3})\neq(0,0,0)$. If $(n_{1},n_{2})=0$, then $\lvert n_{3}\rvert\geq 2$, so (4.2) gives
@@
(4.12) $f(n_{1},n_{2},n_{3})\geq 8\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}.$
@@
Now consider the case $(n_{1},n_{2})\neq(0,0)$. Since $\operatorname{LOG}(u_{l}),\operatorname{LOG}(u_{0}),\operatorname{LOG}(\sigma(u_{0}))~{}$ are perpendicular as seen in (4.1), we have
@@
$=W_{1}^{2}+W_{2}^{2}+W_{3}^{2}$
from (4.10). Then Theorem 4.5 gives
$W_{1}^{2}+W_{2}^{2}+W_{3}^{2}\geq 4\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}.$
Under the constraint of above inequality, (4.5), (4.6), and the fact that  $W_{1},W_{2},W_{3}$ are nonzero, an elementary consideration yields
$2W_{1}(2\max\{\lvert W_{2}\rvert,\lvert W_{3}\rvert\}+\lvert W_{2}\rvert+\lvert W_{3}\rvert)>6\sqrt{3}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}.$
@@
$f(n_{1},n_{2},n_{3})>6\sqrt{3}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}.$
Since $8\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}>6\sqrt{3}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}$, by (4.12), the above inequality holds for any $(n_{1},n_{2},n_{3})\neq(0,0,0)$ with $n_{1}+n_{2}+n_{3}$ even. Therefore, with (4.11),
$\lVert w\rVert_{1}>3\sqrt{3}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}$

### v2

@@ 4. The Cyclic Group Case
In this section, we prove the main theorem for the case when  $\operatorname{Gal}(L/\mathbb{Q})$ is isomorphic to the cyclic group. Throughout this section, let $L$ be a real Galois extension of $\mathbb{Q}$ with $\operatorname{Gal}(L/\mathbb{Q})\cong\mathbb{Z}/4\mathbb{Z}$. Denote the Galois group as  $G$. Let $\operatorname{Gal}(L/\mathbb{Q})$ be generated by $\sigma$. Let $l$ be the unique quadratic subfield of $L$. Let $u_{l}>1$ be the fundamental unit of $l$. Then ${\mathcal{O}_{l}^{*}}=\{\pm u_{l}^{m}\ |\ m\in\mathbb{Z}\}$. First we define relative units as in [5].
@@
Let $E_{L/l}$ be a group consisting of all relative units of $L$. As in [4], let $E^{L}$ be a sub $G$-module of $\mathcal{O}_{L}^{*}$} generated by elements of  $\mathcal{O}_{l}^{*}$} and $E_{L/l}$. Let $Q=[{\mathcal{O}_{L}^{*}}:E^{L}]$. The following two lemmas are due to Hasse [5, Satz 16 and 22].
@@
There exists an element $u_{0}$ of $E_{L/l}$ such that $1,u_{0}$ and its conjugate $\sigma(u_{0})$ generates $E_{L/l}$.
@@
[…]
@@
(4.2) $f(n_{1},n_{2},n_{3})\geq 4\lvert n_{3}Y_{1}\rvert=4\lvert n_{3}\rvert(W_{2}^{2}+W_{3}^{2})$
@@
[…]
Using these results, we prove Theorem 1.2 for each case $Q=1$ and $Q=2$.
@@
[…]
$\lvert mX+nY\rvert+\lvert nX-mY\rvert\geq\lvert X\rvert+\lvert Y\rvert$.
@@
[…]
@@
[…]
$=\lvert m(1+Z^{2})\rvert\;\geq m\;\geq 2\;\geq 1+Z.$
@@
$\lvert 1+nZ\rvert+\lvert n-Z\rvert\geq{-n+Z}\geq 1+Z,$
@@ 4.1. The Case Q=1
[…]
@@
[…]
@@
Let $L$ be a totally real number field and $u\in{\mathcal{O}_{L}^{*}}$ with $u\neq\pm 1$. Then we have
$\lVert\operatorname{LOG}(u)\rVert_{2}^{2}\geq[L:\mathbb{Q}]\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right).$
@@
Bemerkung in p. 102 of Pohst [8] states that the Lemma in p. 98 applies for  $M$ of the equation (10) in p. 97. Then the inequality (11) of the Lemma gives the required result. ∎
@@
(4.5) $W_{2}^{2}+W_{3}^{2}=\frac{1}{2}\lVert\operatorname{LOG}(u_{0})\rVert_{2}^{2}\geq 2\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right).$
Under this constraint, the following elementary fact will be useful.
Lemma 4.6.
Let $x,y\in\mathbb{R}$ with $x^{2}+y^{2}\geq R^{2}$ for some positive real number $R$. Then,
$2\max\{\lvert x\rvert,\lvert y\rvert\}+\lvert x\rvert+\lvert y\rvert\geq 2\sqrt{2}R.$
Proof.
Without loss of generality, assume $x\geq y\geq 0$ and $x^{2}+y^{2}=R^{2}$. Then there exists $\theta\in[0,\tfrac{\pi}{4}]$ such that $x=R\cos\theta$ and $y=R\sin\theta$. Hence
$2\max\{\lvert x\rvert,\lvert y\rvert\}+\lvert x\rvert+\lvert y\rvert=3x+y=\sqrt{10}R\sin(\theta+\phi),$
where $\phi=\arctan(3)$. Since $\theta+\phi$ ranges over $\arctan(3),\arctan(3)+\tfrac{\pi}{4}]\subset(0,\pi)$ and $\sin(\phi)=\tfrac{3}{\sqrt{10}}$ is larger than $\sin(\phi+\tfrac{\pi}{4})=\tfrac{2}{\sqrt{5}}$, the minimum of $\sin(\theta+\phi)$ occurs at $\theta=\tfrac{\pi}{4}$. The minimum value is therefore $\sqrt{10}\,R\cdot\tfrac{2}{\sqrt{5}}\;=\;2\sqrt{2}\,R$. ∎
Combining Lemma 4.6 with (4.5), we obtain
$2\max\{\lvert W_{2}\rvert,\lvert W_{3}\rvert\}+\lvert W_{2}\rvert+\lvert W_{3}\rvert\geq 4\log\left(\tfrac{1+\sqrt{5}}{2}\right).$
@@
(4.6) $W_{1}=\log(u_{l})\geq\log\left(\tfrac{1+\sqrt{5}}{2}\right)$
holds, and with the previous inequality,
$2W_{1}(2\max\{\lvert W_{2}\rvert,\lvert W_{3}\rvert\}+\lvert W_{2}\rvert+\lvert W_{3}\rvert)\geq 8\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right).$
@@
(4.7) $\lVert w\rVert_{1}\geq 8\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right)$
@@ 4.2. The Case Q=2
We use the following lemma due to Hasse [5, Satz 28].
@@
If $Q=2$, there exists a unique unit $u_{*}>0$ of $L$ such that
$|N_{L/l}(u_{*})|=u_{l},\qquad|u_{*}\sigma(u_{*})|=\lvert u_{0}\rvert$.
@@
and therefore  $\sigma(u_{0})=\pm u_{l}u_{0}u_{*}^{-2}$. The proof of Lemma 4.2 shows that when $Q=2$, every unit $u\in{\mathcal{O}_{L}^{*}}$ can be written uniquely as the form $\pm u_{*}^{n_{0}}u_{l}^{n_{1}}u_{0}^{n_{2}}\sigma(u_{0})^{n_{3}}$ where $n_{0}$ is 0 or $1$ and $n_{1},n_{2},n_{3}$ are integers. Thus  $\mathcal{O}_{L}^{*}$} is generated by $1,u_{l},u_{0},u_{*}$. Hence  $\bigwedge^{2}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$ is generated by $\operatorname{LOG}(u_{l})\wedge\operatorname{LOG}(u_{0}),\operatorname{LOG}(u_{l})\wedge\operatorname{LOG}(u_{*})$ and  $\operatorname{LOG}(u_{0})\wedge\operatorname{LOG}(u_{*})$. From (4.8), we also have
@@
We now proceed to give a lower bound for $f(n_{1},n_{2},n_{3})$ where $n_{1}+n_{2}+n_{3}$ is even and $(n_{1},n_{2},n_{3})\neq(0,0,0)$. If $(n_{1},n_{2})=(0,0)$, then $\lvert n_{3}\rvert\geq 2$, so (4.2) gives
@@
(4.11) $f(n_{1},n_{2},n_{3})\geq 16\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right).$
@@
Now consider the case $(n_{1},n_{2})\neq(0,0)$. Since $\operatorname{LOG}(u_{l}),\operatorname{LOG}(u_{0}),\operatorname{LOG}(\sigma(u_{0}))$~ are perpendicular as follows from (4.1), we have from (4.9)
@@
$=W_{1}^{2}+W_{2}^{2}+W_{3}^{2}.$
Then Theorem 4.5 gives
$W_{1}^{2}+W_{2}^{2}+W_{3}^{2}\geq 4\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right).$
Letting $S=\sqrt{W_{2}^{2}+W_{3}^{2}}$, the inequality becomes
(4.12) $W_{1}^{2}+S^{2}\geq 4\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right),$
while (4.5) gives $S\geq\sqrt{2}\log\left(\tfrac{1+\sqrt{5}}{2}\right)$.
We claim that
(4.13) $W_{1}S\geq\sqrt{3}\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right).$
Indeed, if $W_{1}\geq\sqrt{2}\log\left(\tfrac{1+\sqrt{5}}{2}\right)$, then since $S\geq\sqrt{2}\log\left(\tfrac{1+\sqrt{5}}{2}\right)$,
$W_{1}S\geq 2\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right)>\sqrt{3}\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right).$
Otherwise, by(4.6), we have $W_{1}\in\left[\log\left(\tfrac{1+\sqrt{5}}{2}\right),\sqrt{2}\log\left(\tfrac{1+\sqrt{5}}{2}\right)\right)$. Then
$W_{1}S\geq\sqrt{W_{1}^{2}(4\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right)-W_{1}^{2})}=\sqrt{4\log^{4}\left(\tfrac{1+\sqrt{5}}{2}\right)-\left(W_{1}^{2}-2\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right)\right)^{2}}.$
This expression attains its minimum at $W_{1}=\log\left(\tfrac{1+\sqrt{5}}{2}\right)$, yielding $\sqrt{3}\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right)$. This proves the claim.
Combining Lemma  4.6 with (4.13), we obtain
$2W_{1}(2\max\{\lvert W_{2}\rvert,\lvert W_{3}\rvert\}+\lvert W_{2}\rvert+\lvert W_{3}\rvert)\geq 4\sqrt{2}W_{1}S\geq 4\sqrt{6}\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right).$
@@
$f(n_{1},n_{2},n_{3})\geq 4\sqrt{6}\log\left(\tfrac{1+\sqrt{5}}{2}\right)^{2}.$
Since $16\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right)>4\sqrt{6}\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right)$, by (4.11), the above inequality holds for any $(n_{1},n_{2},n_{3})\neq(0,0,0)$ with $n_{1}+n_{2}+n_{3}$ even. Therefore, with (4.10),
$\lVert w\rVert_{1}\geq 2\sqrt{6}\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right)$